OpenStudy (anonymous):

WILL MEDAL AND FAN!!!What is the measure of Z? A. 76° B. 82° C. 101° D. 152°

OpenStudy (anonymous):

@KamiBug

OpenStudy (bigdaddyharsha):

A

OpenStudy (bigdaddyharsha):

A) 76

jimthompson5910 (jim_thompson5910):

The three red curved bands for angles Y and Z tell us that angle Y is congruent to angle Z So you can say Y = Z

jimthompson5910 (jim_thompson5910):

`W+X+Y+Z = 360` because for ANY quadrilateral (figure with 4 sides), the 4 angles always add to 360 degrees

jimthompson5910 (jim_thompson5910):

so, W+X+Y+Z = 360 129+79+Y+Z = 360 ... plug in the values for X & W 129+79+Z+Z = 360 ... replace Y with Z (since Y = Z) now solve for Z